# An End-to-End Room Geometry Constrained Depth Estimation Framework for Indoor Panorama Images

### v1

@@ IV-B1 Quantification Comparison with SoTA Depth Estimation Methods
On the Structured3d dataset, we mainly compare with BGDNet, which uses a similar idea. The results released in the BGDNet paper are trained on the replica and then val on the structure3d dataset, while our method is trained on Structured3d and then validated on the corresponding validation set. This comparison may not be fair, but considering that the BGDNet paper is not open source, we can only use this method for comparison. Due to the differences in the settings of train and val, the method proposed in this paper has a very obvious advantage in the six indicators of concern.
TABLE IV: Quatitative analysis of our proposed dataset denoise method.
Dataset Setting Methods RMSE MRE MAE
Without Denoise PanoFormer 0.3083 0.0405 0.1013
Our Methods 0.2673 0.0375 0.0912
Denoised Dataset PanoFormer 0.2872 0.0392 0.0892
Our Methods 0.2359 0.0285 0.0580
TABLE V: The training impact of each task decoder to depth decoder
@@
0.3635 0.0571 0.1013
0.3621 0.0531 0.1011
0.3656 0.0582 0.1101
0.3592 0.0511 0.1002
TABLE VI: The accuracy of resolved background depth maps.
Methods RMSE MRE MAE
PanoFormer’s Background Region 0.3214 0.0623 0.1123
Decode from layout prediction 0.0043 0.0032 0.0023
Decode from layout ground-truth 0.0001 0.0001 0.0001
TABLE VII: The comparison of our proposed fusion methods and BGDNet
Methods RMSE MRE MAE
Baseline 0.3635 0.0571 0.1013
Fusion BGDNet 0.2823 0.0498 0.0922
Fusion ours 0.2436 0.0425 0.0829
@@
Figure 7: The experiment results of explore the inference of hyper-parameter $\lambda_{2}$ and $\lambda_{3}$. The x-axis in this figure represents the value of the corresponding hyper-parameter setting, while y-axis represents the RMSE value of our proposed framework’s depth prediction.
IV-B2 Visualize Comparison with SoTA Depth Estimation Method
In this section, we perform visualization operations on the scene as shown in the figure. In this set of experiments, we selected the baseline panoformer in this paper as the comparison object. It is worth noting that in this set of experiments, panoformer and the depth estimation framework proposed in this paper both use denoised datasets during training. In this set of experiments, we selected several representative scenes on the validation set for demonstration. From the corresponding 3D visualization results, it can be seen that the depth estimated by our method is significantly more accurate in corners and walls. At the same time, thanks to the prior knowledge of room layout introduced by us, the 3D visualization of the prediction results of our method is more reasonable in terms of overall structure. At the same time, due to the existence of our dataset denoising strategy, panoformer and our method do not have the phenomenon of overfitting of noise points in patches as shown in Fig. 1. The existence of these phenomena fully demonstrates the effectiveness of the depth estimation method we proposed.
IV-C Visualize and Quatitative analysis of Dataset Denoise Method
IV-C1 Visualize Analysis of Denoise Method
We conducted a visualization experiment as shown above for the panoramic depth images of the 2D-3D-S dataset collected in the real world. After mapping each pixel in the panoramic depth images to 3D space, we found that there are a large number of noise points in each scene. For these noise points, the processing strategy of the existing panoramic depth estimation method is to directly preset a maximum distance threshold. In the 2D-3D-S dataset, this threshold is set to 10 meters. The depth values of all areas greater than this threshold will be artificially set to 0. On the one hand, this processing method destroys the correlation between the pixels of the original panoramic depth map, thereby affecting the learning process of the model. On the other hand, it will also cause some areas in the scene with depth values less than 10 but should fall inside the room to fall outside the room. In order to preserve the pixel correlation of the processed panoramic depth map and prevent the area falling inside the room from falling outside the room, we proposed a strategy to preprocess the indoor panoramic depth map using pretrained layout. In this strategy, we first collected as many parts of the existing dataset that contain layout annotations as possible to form a layout dataset of a relatively reasonable size to pretrain the layout estimation part of the framework proposed in this paper. Considering that the layout prediction branch can already obtain a high 3D IoU in our experiments, we directly use this part of the prediction results to process the depth estimation dataset in the actual processing process. In our proposed processing strategy, we first use the layout prediction branch to estimate the layout of the current room, and then use the panorama depth ground-truth to count the height of each wall of the room, and use this height result to solve the 2D layout into the corresponding background depth map. Finally, we use this background depth map to constrain the original panorama depth ground-truth as shown in Fig. 6. In this constrained process, we use the background ground-truth as a mask to ensure that the depth of the background region does not exceed the range of the room itself.
IV-C2 Quatitative Analysis of Denoise Method
We use the Stanfor2d3d dataset as the object to verify the effectiveness of the proposed noisy image processing strategy. In the experiment, we use our proposed framework and PanoFormer to perform training and val on the processed datasets before and after processing. The experimental results are shown in Table IV. From the results in Table IV, the performance of PanoFormer and our proposed framework trained on the processed dataset is significantly higher than that trained on the original dataset, which fully demonstrates the performance improvement brought by our data processing strategy.
IV-D Ablation Study
IV-D1 The Effect of Framework’s Each Component
Figure 8: The visualize of Background segmentation. The left column of the figure is input panorama RGB images, the middle column is background segmentation predictions, the right column is corresponding ground-truth.
Figure 9: The visualize of room layout estimation. The first column shows the input RGB image. The second column shows the room layout estimation results displayed on a 2D equirectangular image. The top row in each 2D equirectangular sub-image represents the probability of a corner point in that column. The third and fourth columns show the corresponding 3D bird’s-eye views of the prediction and layout ground-truth, respectively.
@@ IV-D1 The Effect of Framework’s Each Component
In this section, we analyze the impact of each proposed component on the whole. Considering that we proposed a framework for auxiliary-task assisted depth estimation tasks, we mainly discuss two aspects of the impact of each component: 1) The impact of auxiliary-task decoder for depth estimation decoder; 2) The effect of our proposed background depth resolving and fusion strategy. The experimental dataset we chose is Matterport3D because it contains complete layout estimation annotations.
Regarding the aspect 1), the corresponding experimental results have been put into Table V. It is worth noting that the loss weight we used in this group of experiments is the combination verified in subsequent section. From the results, it can be seen that the auxiliary-task decoder has very little impact on the depth decoder during training. This is primarily because we intentionally set a low loss weight for the auxiliary-task decoder. This design stems from two considerations: first, optimizing each task in a multi-task learning framework requires extensive experimentation and manual tuning. Second, and more importantly, our framework’s primary objective is to use the auxiliary task’s output to constrain the depth decoder directly, rather than to improve it indirectly by learning more generalized features in the shared encoder.
Regarding the aspect 2) we want to discuss, we conducted corresponding experiments and put them in Table VI and Table VII. The experimental results of the background depth resolving strategy are shown in Table VI, and the experimental results of the fusion strategy used are shown in Table VII. When evaluating the background depth resolving strategy, we used the depth ground-truth provided by the Matterport3D dataset as a benchmark. When evaluating the resolved background depth, we used the ground-truth of the background segmentation as a mask to ensure that only the background region is evaluated during the eval process. Judging from the results in Table VI, our background depth resolving strategy is more accurate than the deep learning model. In terms of fusion strategy, we chose BGDNet[32] as the comparison object. In this set of experiments, the configuration of the fusion strategy used by BGDNet is consistent with the original paper. The experimental results demonstrate that incorporating background depth as a constraint effectively enhances depth estimation performance. Moreover, our proposed fusion strategy yields even more substantial improvements compared to that of BGDNet.
@@
IV-D2 Analysis of loss weights
The impact of different loss weights hyper-parameters of the proposed framework has been shown in Fig. 7. Here, we use the MatterPort3D to conduct experiments. As the layout decoder in our framework has been pretrained, we simply fix its $\lambda_{1}$ to 0.01. For $\lambda_{2}$ and $\lambda_{3}$, we first fix the $\lambda_{3}$ to 1.0 and search the best setting of $\lambda_{2}$. Then, fix the $\lambda_{2}$ to the searched results and search the setting of $\lambda_{3}$. The Fig. 6. show that on the MatterPort3D dataset, when we set $\lambda_{2}=1.0$ and $\lambda_{3}=0.4$, the proposed framework can achieve the best performance.
IV-E Visualize Analysis of Background Segmentation and Layout Estimation
In this section, we visualize the prediction results of the front-segmentation and layout estimation task. We used the test set data of Structured3D to perform correspond experiment. The specific visualization results are shown in Fig. 8 and Fig. 9.
Fig. 8 shows the visualization results of the background segmentation task. In our framework, the prediction results of the background segmentation task are used as weight to fuse the coarse prediction and background depth. When visualizing the results of background segmentation, we mapping the prediction score maps from $[0,1]$ to $[0,255]$ and save them as gray images. Judging from the score maps shown in Fig. 8, our background branch can accurately distinguish between the background and background areas in the input image. However, in some areas that are very close to the wall and have a similar depth to the background area, the discrimination ability of our background branch still needs to be improved.
Fig. 9 shows the estimation results of our layout estimation branch. From the 2D visualization of the prediction results, we can see that the edge areas and corner points of the wall can be identified relatively accurately. However, in some areas far away from the camera, the prediction accuracy of the corner points still needs to be improved. In addition to these issues, another point worth noting is that the current room layout estimator assumes that the walls of the room are all cubic structures. This makes our method perform worse than the normal depth estimator when estimating the depth of some rooms with curved wall structures. We consider conducting more detailed research on this issue in future work.
V Conclusions
This paper has proposed a panoramic depth estimation framework with room geometry constraints. The framework employs multi-task learning, where a shared encoder extracts features that are decoded into three outputs: coarse depth maps, background segmentation masks, and room layouts. These predictions enable our method to initially reconstruct background depth from the layout information. Subsequently, background depth and coarse depth are fused using the background segmentation mask as a weighting mechanism, ultimately generating the final depth prediction. Extensive experiments on real-world and synthetic datasets demonstrate significant performance improvements over current methods.

### v2

@@ IV-C1 Quantification Comparison with SoTA Depth Estimation Methods
On the Structured3d dataset, we mainly compare with BGDNet, which uses a similar idea. The results released in the BGDNet paper are trained on the replica and then val on the structure3d dataset, while our method is trained on Structured3D and then validated on the corresponding validation set. Due to the differences in the settings of train and val, the method proposed in this paper has a very obvious advantage in the six indicators of concern.
On the Replica dataset, we ensured that all data processing steps were as consistent as possible with BGDNet. Based on the performance across the six metrics shown, our method significantly outperforms BGDNet. Furthermore, it demonstrates a clear advantage over the current state-of-the-art methods, DepthAnyPanorama and DepthAnyDirection. It’s worth noting that, considering both of these methods used large datasets to pre-train large models, the performance we present is the result of fine-tuning the officially provided weights on the Replica dataset for two epochs.
Overall, in terms of quantitative metrics, our method demonstrates a significant performance advantage compared to current state-of-the-art methods, especially in the most common metric rmse. This clearly shows that using the background depth we constructed effectively constrains the depth estimator’s predictions, leading to more accurate depth estimation results.
IV-C2 Visualize Comparison with SoTA Depth Estimation Method
In this section, we provide a qualitative evaluation of depth estimation results, with visual comparisons shown in Fig. 4. The selected comparison methods include HoHoNet[29], a multi-task learning approach, and EGFormer[38], which is based on a panoramic Transformer architecture. The visual results demonstrate that, owing to the effective constraint provided by our background depth modeling, our framework captures the overall geometric structure of the room more accurately than existing open-source state-of-the-art methods. This improvement also validates the accuracy of the enclosed-region background depth recovered by our pose-aware depth resolving strategy.
It should be noted that while our framework predicts the overall room geometry more accurately, its performance remains limited in areas outside the enclosed background region. Specifically, in the first scene, distant objects such as the sofa are not well reconstructed. In the second scene, certain protruding structures are incorrectly compressed onto the wall planes in the depth prediction.
The limitation observed in the first scenario is primarily attributed to the fact that the current enclosed-region layout estimation can only represent a limited portion of the overall room structure. For depth prediction in areas beyond the enclosed region, our framework relies principally on the capacity of the base depth decoder. Nevertheless, accurate depth estimation in such areas remains a persistent challenge for panoramic depth estimators in indoor scenes, and a widely effective solution has yet to be established.
The issue observed in the second scenario can be attributed to current limitations in irregular region segmentation. Inherent inaccuracies in this task lead to certain areas outside the enclosed background being misclassified as background regions. Consequently, these areas—which should reflect the room’s extended structure—are incorrectly flattened onto the wall planes in the depth prediction.
IV-D Quantification Analysis of Background Depth Resolving
TABLE IV: The performance of our framework’s layout decoder
Methods 2DIoU 3DIoU RMSE $\delta^{1}$
HorizonNet[28] 91.18 89.77 0.09 0.99
HoHoNet[29] 92.96 91.52 0.09 0.99
Ours 94.36 92.96 0.04 0.99
TABLE V: The accuracy of our framework’s pose decoder
Methods RRA@30 RTA@30 AUC@30
Ours 94.33 90.51 91.34
TABLE VI: The accuracy of camera height
mean average error
camera height from pose decoder $\tilde{h}_{c}$ 0.05
camera height from depth decoder $\hat{h}_{c}$ 0.04
camera height from resolving strategy $h_{c}$ 0.02
TABLE VII: The accuracy of resolved background depth maps.
Methods RMSE MRE MAE
HoHoNet’s Background 0.3514 0.1223 0.1923
PanoFormer’s Background 0.2644 0.0688 0.0991
Layout GT with Fix Camera pose 0.5514. 0.3846 0.3102
Layout GT with GT Camera pose 0.0011 0.0010 0.0010
Layout GT with Pred Camera pose 0.0013 0.0012 0.0012
Layout Pred with Fix Camera pose 0.5862 0.3212 0.3833
Layout Pred with GT Camera pose 0.0403 0.0312 0.0213
Layout Pred with Pred Camera pose 0.0551 0.0417 0.0412
In this section, we will construct corresponding experiments to verify and compare the performance differences between our background depth calculation strategy and BGDNet’s background depth calculation strategy. For convenience, we will only construct experiments on the Structured3D dataset. We will first demonstrate the overall performance of the framework’s layout decoder and camera pose parts. Based on this, we will show the RMSE, MRE, and MAE of the background region depth and depth ground truth calculated by layout ground truth and layout prediction with and without camera pose. Furthermore, we will demonstrate the corresponding metrics of the background depth calculated using our method based on camera pose.
The results of layout estimation and camera pose estimation in the framework are shown in Tables IV and V, respectively. Table VI compares the camera height $\hat{h}_{c}$ calculated from the depth decoder, the camera height $\tilde{h}_{c}$ obtained from the camera pose decoder, and the average of the two $h_{c}$ in our pose-aware background depth resolving strategy. The results in Tables IV and V show that the layout decoder and pose estimation results in our framework are quite accurate. However, the results in Table 6 show that the camera height calculated from either the camera pose decoder or the depth decoder still has an error of about 5 cm; however, the error of the camera height obtained by averaging the two is reduced to 2 cm. This fully demonstrates the effectiveness and rationality of the pose estimation strategy in our proposed method.
Table VII presents the accuracy of our baseline model in background depth and the experimental results related to background depth calculated based on layout. The results show that the depth prediction results of our chosen open-source baseline model are relatively poor in the background portion. Furthermore, the background depth value calculated by room layout is highly dependent on the camera height. The results in rows 3 and 6 indicate that the camera pose used in generating images in different virtual scenes varies in the structured3d dataset. Rows 4 and 7 show that the background depth generated by ground truth layout can be considered, to some extent, as ground truth depth. The experimental results in rows 5 and 8 show that the background depth obtained by our proposed background depth calculation strategy also has some minor errors, but these errors are still orders of magnitude different from those produced by other current depth estimators.
IV-E Quantification and Visualize Analysis of Fusion Mask Generation Component
TABLE VIII: The mIoU of background segmentation of our region segmentation decoder
Methods mIoU
HoHoNet[29] 72.12
SAN[36] 95.33
Ours 89.23
TABLE IX: The mIoU of irregular region segmentation of our region segmentation decoder
Methods mIoU
HoHoNet[29] 73.33
SAN[36] 96.32
Ours 90.92
Figure 5: The visualize of resolved background depth. The figure shows 3D visualizations of the background depth in four scenes constructed using our method and BGDNet. The areas marked with red circles represent regions that BGDNet misjudged as exceeding the background depth, while the areas marked with yellow circles represent regions that our constructed background depth accurately identifies.
In this section, we construct experiments for the region segmentation decoder and the Fusion Mask Generation Component and verify their performance. The dataset used in the experiment is the same as that in the previous subsection, using Structured3D. Tables VIII and IX show the performance of the two semantic segmentations performed by the region segmentation decoder. In these two subtasks, we chose the open-source HoHoNet as the comparison object. We did not choose newer methods because recent methods using multi-task learning on panoramic images are not open-source, so we could not conduct corresponding experiments on our own constructed experimental setup. Regarding our Fusion Mask Generation Component, the visualization results of our constructed experiments are shown in Fig. 5.
The experimental results in the two tables show that HoHonet exhibits significantly higher performance than the original paper due to the substantial reduction in task difficulty. However, considering HoHonet’s overly aggressive approach of compressing 2D features of panoramic images to 1D, its performance still lags considerably behind our method. It’s worth noting that while our method represents a significant performance improvement over earlier multi-task learning methods on panorama, it still lags behind recent methods specifically designed for semantic segmentation tasks, such as SAN[36]. We believe this difference largely stems from our reuse of a portion of the depth decoder network for model overhead considerations.
As shown in Fig. 5, BGDNet does not consider that the depth of some regions in panoramic images in real-world scenes may exceed the space defined by the room layout when constructing the background depth. This leads to a disastrous problem for BGDNet when constructing the background depth: regions that should extend beyond the background depth are compressed onto the walls defined by the room layout. Although the binary semantic segmentation mask provided by SAM used by BGDNet can effectively distinguish which regions are walls, floors, and ceilings, the semantic segmentation mask provided by SAM cannot help the model identify which regions extend beyond the room layout, and these regions also belong to walls, floors, or ceilings. In contrast, our method benefits from the irregular region mask provided by the region segmentation decoder, which can effectively distinguish which parts of the scene extend beyond the room layout. It is worth noting that since the irregular region mask and background mask themselves are not completely accurate, the background depth constructed by our method has fewer regions for fusion, which reduces the constraint effect of the background depth to some extent. However, we believe that reducing some of the blended areas to make the background depth itself more generalizable is a reasonable trade-off.
IV-F Quantification Analysis of Adaptive Fusion Component
TABLE X: The comparison of our proposed fusion methods and BGDNet
Methods RMSE MRE MAE
Baseline 0.3017 0.1201 0.1546
Fusion BGDNet[6] 0.2643 0.0812 0.1122
Fusion ours 0.1953 0.0414 0.0613
In this section, we construct corresponding experiments to verify the performance of our Adaptive Fusion Component. The relevant experimental results are shown in Table X. The dataset used in the experiments remains structured3d. It is worth noting that, to ensure a fair comparison, we also used the irregular region mask provided by the region segmentation decoder of our framework in the experiments against BGDNet to ensure the accuracy of its background depth. As can be seen from the experimental results in Table X, our fusion strategy is significantly better than BGDNet’s approach. From the perspective of model structure, the region segmentation decoder in our framework is functionally equivalent to the module in BGDNet that extracts and fuses background depth features, while the subsequent threshold replacement operation in BGDNet functionally corresponds to our adaptive fusion component. Based on these analyses, we speculate that the advantage of our fusion method stems from our explicit learning of the weight process for distinguishing foreground and background regions as a supervised learning task during training. This learning-based approach is superior to implicitly fusing information into the feature extraction network.
IV-G Ablation Study
TABLE XI: The training impact of each task decoder to depth decoder
@@ IV-G Ablation Study
0.3017 0.1201 0.1546
0.3011 0.1109 0.1411
0.2993 0.1182 0.1424
0.3012 0.1192 0.1504
TABLE XII: The training impact of each component to the final depth predictions.
Background
Resolving
Fusion Mask
Generation
Adaptive
Fusion
RMSE MRE MAE
0.3017 0.1201 0.1546
0.2843 0.1011 0.1322
0.2231 0.0672 0.0833
0.1935 0.0414 0.0613
@@
Figure 6: The experiment results of explore the inference of hyper-parameter $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$. The x-axis in this figure represents the value of the corresponding hyper-parameter setting, while y-axis represents the RMSE value of our proposed framework’s depth prediction.
@@
In this section, we analyze the impact of each proposed component on the whole. Considering that we proposed a framework for auxiliary-task assisted depth estimation tasks, we mainly discuss two aspects of the impact of each component: 1) The impact of hyper-parameter setting for final depth predictions ;2) The impact of auxiliary-task decoder for depth estimation decoder; 3) The impact of each component for the final depth predictions. The experimental dataset we chose is Structured3D because it contains complete layout estimation annotations.
Regarding aspect 1) which we want to discuss, the relevant experimental results are shown in Fig. 6. In this set experiments, we divided the weight values of each component in the objective function within the range 0-1 into 0-1 into intervals of 0.1. We adopted a component-by-component exploration mode for configuring the optimal loss weights. Each time we explored the optimal configuration for a parameter, we set all subsequent parameters to 0 while keeping the preceding parameters unchanged. Considering that the layout decoder and other task decoders are trained independently in our framework, the value of its corresponding parameter $\lambda_{1}$ is constant at 0 when training other decoders. As shown in Fig. 6, the framework achieves optimal overall performance when $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are set to 1.0, 0.4, and 0.6, respectively.
Regarding aspect 2), the relevant experimental results are shown in Table XI. In this set of experiments, we simply discuss the impact of each task decoder involved in the training process on the depth decoder results. In this set of experiments, the Loss weights for each task decoder used the optimal configuration explored in the previous set of experiments. As the results shown in the table, different task decoders do not have a significant impact on the final result during training in our framework. This also suggests that the parameters explored in the previous set of experiments actually have a greater impact on the final experimental results than on the performance of each task decoder, which in turn affects the performance of the three components we proposed, ultimately impacting the overall framework performance.
Regarding aspect 3), the relevant experimental results are shown in Table XII. In this set of experiments, we used the optimal parameters explored for each task decoder. The results in the first row of Table XII. represent the experimental results of our baseline Panoformer; the second row represents the performance when we used the proposed background depth resolving component in conjunction with BGDNet’s background replacement strategy; and the third and fourth rows represent the performance when we combined our proposed fusion mask generation and adaptive fusion components, respectively. From the results in Table XII., it is clear that the core of our framework’s performance improvement lies in the three task components rather than multi-task learning, and the fusion mask generation component shows the most significant improvement. This experimental result also fully validates the core argument of this paper: to make the background depth constraint more reasonable, we need to more carefully identify which regions in the input indoor panorama image belong to the background.
@@
V Limitation
The proposed PAGCNet achieves significantly superior performance on current indoor depth estimation datasets. However, our framework still has several issues:
1) To adapt to more complex indoor scenes, we divide the scene into regular and irregular regions and only model the background for the regular regions. Therefore, we don’t actually model the irregular regions, and naturally, we don’t apply geometric constraints to their depth. This makes our framework perform poorly on some extremely irregular room structures, although these scenes are very rare.
2) In existing panoramic depth estimation datasets, the number of annotations for the room layout task cannot be matched one-to-one with the annotations for tasks like semantic segmentation and depth estimation. This mismatch forces us to train the room layout decoder separately before training other decoders. Therefore, our framework doesn’t actually solve this problem of annotation imbalance between tasks.
VI Conclusions
This paper has proposed a pose-aware and geometry-constrained framework for irregular panoramic depth estimation.Our framework first employs multiple task-specific decoders to jointly estimate room layout, camera pose, depth, and region segmentation from a single input panorama. A pose-aware background depth resolving component uses tasks decoder’s prediction to refine the camera pose and subsequently uses the camera pose to compute the background depth of regular enclosed regions, which serves as a strong geometric prior. Based on the output of the region segmentation decoder, a fusion mask generation component produces a fusion weight map to guide where and to what extent the geometry-constrained background depth should correct the depth decoder’s prediction. Finally, an adaptive fusion component integrates this refined background depth with the initial depth prediction, guided by the fusion weight. Extensive experiments on Matterport3D, Structured3D, and Replica datasets demonstrate that our method achieves significantly superior performance compared to current open-source methods.